list down all the subsets of E= (1,2,3)
Answer by 2897696(96)

to know how many subsets of E use formula 2^(number in the set)
2^3=2*2*2=8 so there are 8 subsets (you can use this knowledge to check your answers later.)
1. frist find all the sets with one number in each subset:(1) (2) (3)
2. next find all the sets with two in each subset: (1,2) (1,3) (2,3)
3. now find all the sets with three in each subset: (1,2,3)
4. *IMPORTANT no solution is a subset to any set: NO SOLUTION
CHECKING:
Count the number of subsets found to see if you found all of them, which is 8.
Since there are 8 solutions, it is correct.
Answer: (1);(2);(3);(1,2);(1,3);(2,3);(1,2,3); NO SOLUTION